OpenStudy (anonymous):

x^2+2x-15 I really just don't understand.

OpenStudy (anonymous):

it means solve X X'2+2X-15=0 X'2-3X+5X-15=0 x(x-3)+5(X-3)=0 (x-3)=0 0r (x+5)=0 X=3 or X=-5